### v2

# Edge open packing on subclasses of chordal graphs

Abstract
Let $G=(V,E)$ be a graph where $V(G)$ and $E(G)$ are the vertex and edge sets, respectively. In a graph $G$, two edges $e_{1},e_{2}\in E(G)$ are said to have a common edge $e\neq e_{1},e_{2}$ if $e$ joins an endpoint of $e_{1}$ to an endpoint of $e_{2}$ in $G$. A subset $D\subseteq E(G)$ is called an edge open packing set in $G$ if no two edges in $D$ share a common edge in $G$, and the largest size of such a set in $G$ is known as the edge open packing number, represented by $\rho_{e}^{o}(G)$. The Maximum Edge Open Packing Problem is to find an edge open packing set of a given graph with maximum size. In [Brešar and Samadi. Edge open packing: complexity, algorithmic aspects, and bounds. Theor. Comput. Sci., 2024.], Brešar and Samadi pose an open question of the edge open packing problem in chordal graphs. In this paper, we partially answer this open question by showing a polynomial-time algorithm to solve the maximum edge open packing problem in the subclasses of chordal graphs. First, we show that the Maximum Edge Open Packing Problem can be solved in polynomial time for proper interval graphs. Furthermore, we show that in block graphs we can solve this problem in polynomial time. Finally, we prove that this problem can be solved in linear time for split graphs.
Keywords. Edge open packing, Interval graphs, Block graphs, Split graphs, Polynomial-time algorithm.
1 Introduction
In this paper, we consider $G$ as a finite simple graph with vertex set $V(G)$ and edge set $E(G)$. Any terms and notation not explicitly defined here follow the definitions in [west2001introduction]. A graph $H=(V^{\prime},E^{\prime})$ is called a subgraph of a graph $G=(V,E)$ if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. Given a vertex subset $S\subseteq V$, the induced subgraph $G[S]$ is the subgraph with vertex set $S$ and edge set consisting precisely of those edges in $E$ whose endpoints are both in $S$. Similarly, for an edge subset $B\subseteq E(G)$, $G[B]$ denotes the subgraph of $G$ induced by the endpoints of edges in $B$. Let $D\subseteq E(G)$. An edge-induced subgraph on $D$ of $G$ is defined as the subgraph of $G$ whose edges are the edges from $D$ and the vertices are the endpoints of edges of $D$. This subgraph is denoted by $G\langle D\rangle$. The open neighbourhood of a vertex $x\in V(G)$, denoted $N_{G}(x)$, is the set of vertices adjacent to $x$. The closed neighbourhood of $x$, denoted $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup{x}$. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. For a connected graph $G=(V,E)$, the distance between two vertices $u,v\in V$, denoted by $d_{G}(u,v)$, is the length (number of edges) of a shortest path connecting them. The diameter of $G$, written as $\operatorname{diam}(G)$, is the maximum distance between any two vertices in $G$, formally defined as $\operatorname{diam}(G)=\displaystyle\max_{u,v\in V}d_{G}(u,v)$. A graph is called bipartite if its vertex set can be divided into two disjoint independent sets. A star is the complete bipartite graph $K_{1,t}$ for $t\geq 1$.
An induced matching $M$ of a graph $G$ is a subset $M\subseteq E(G)$ such that $M$ constitutes a matching and no two edges in $M$ are connected by any edge of $G$. The induced matching number $G$ represents the maximum size of an induced matching within $G$. The problem of finding a maximum induced matching was first presented by Stockmeyer and Vazirani as the “risk-free marriage problem” [stockmeyer1982np] and has since been thoroughly examined (see, for instance, [cameron1989induced, dabrowski2013new, lozin2002maximum]).
In 2022, Chelladurai et al. [chelladurai2022edge] introduced the concept of an edge open packing (EOP) in graphs. A subset $D\subseteq E(G)$ is termed an edge open packing set (EOP set) if, for every pair of distinct edges $e_{1},e_{2}\in D$, no edge in $E(G)\setminus{e_{1},e_{2}}$ connects endpoints of $e_{1}$ with endpoints of $e_{2}$. Equivalently, for any two edges $e_{1},e_{2}\in D$, there is no third edge $e_{3}\in E(G)\setminus{e_{1},e_{2}}$ forming either a path $P_{4}$ (with four vertices) or a cycle $C_{3}$ (with three vertices) in the subgraph $G\langle e_{1},e_{2},e_{3}\rangle$. The edge open packing number, denoted by $\rho^{o}_{e}(G)$, is defined as the maximum size of an EOP set in $G$. An EOP set with cardinality equal to $\rho^{o}_{e}(G)$ is called a $\rho^{o}_{e}$-set. It is noteworthy that if $D$ is an EOP set (resp. an induced matching), the induced subgraph $G[D]$ is a disjoint union of induced stars (resp. $K_{1,1}$-stars). Thus, the induced matching problem can be viewed as a special case of the EOP problem, which relaxes the stricter conditions of a matching or induced matching.
An injective $k$-edge coloring of a graph $G$ is an assignment of colors (integers) from the set $\{1,\ldots,k\}$ to the edges of $G$ such that any two edges $e_{1},e_{2}\in E(G)$ that share a common adjacent edge, meaning an edge $e$ exists with one endpoint common to $e_{1}$ and the other endpoint common to $e_{2}$, are assigned distinct colors. The smallest integer $k$ for which this coloring exists is known as the injective chromatic index of $G$, denoted by $\chi^{\prime}_{i}(G)$. This notion was introduced by Cardoso et al. [cardoso2019injective] and subsequently explored in greater depth in [foucaud2021complexity, miao2022note]. It is noteworthy that each color class generated by an injective edge coloring constitutes an edge open packing (EOP) set. Consequently, determining $\chi^{\prime}_{i}(G)$ is effectively the problem of partitioning the edge set $E(G)$ into EOP sets. Practical applications of injective edge colorings and edge open packings have been discussed in the literature, specifically in [cardoso2019injective] and [chelladurai2022edge].
In [chelladurai2022edge], Chelladurai et al. studied edge open packing in graphs. They showed some bounds in terms of a graph’s diameter, size, minimum degree, clique number, and girth. Also, they characterized graphs $G$ with $\rho_{e}^{o}(G)\in\{1,2,m-2,m-1,m\}$, where $m$ is the number of edges of $G$. Further, Brešar and Samadi [brevsar2024edge] study the EOP set problem from an algorithmic point of view. They proved that the decision problem associated with Maximum Edge Open Packing Problem is NP-complete for three special families of graphs, namely, graphs with universal vertices, Eulerian bipartite graphs, and planar graphs with maximum degree $4$. Moreover, they designed a linear-time algorithm in trees to solve the Maximum Edge Open Packing Problem. Also, they provided a structural characterization of all graphs $G$ attaining the upper bound $\rho_{e}^{o}(G)\leq|E(G)|/\delta(G)$. They analyzed the effect of edge removal on the EOP number of a graph $G$, obtaining a lower and an upper bound for $\rho_{e}^{o}(G-e)$, where $e$ is an edge of $G$. In [BRESAR2025], Brešar et al. showed some relationship between the induced matching number and the EOP number in trees and graph products. Later, Pandey and Santra [pandey2025edge] further characterize this parameter for any $t\geq 3$. They showed the necessary and sufficient condition for $\rho_{e}^{o}=t$, for $t\geq 3$, and explicitly defined graph families for $\rho_{e}^{o}=m-3$, where $m$ is the number of edges of $G$. In[brevsar2024edge], Brešar and Samadi pose the following open question of the edge open packing problem in chordal graphs.
Question 1.
Is there an efficient algorithm to determine the edge open packing number of a chordal graph ??
In this paper, we partially answer this open question by showing a polynomial-time algorithm to solve the maximum edge open packing problem in the subclasses of chordal graphs, namely proper interval graphs, block graphs and split graphs.
An edge between two non-consecutive vertices of a cycle is called a chord. If every cycle in $G$ of length at least four has a chord, then $G$ is called a chordal graph. A vertex $v\in V$ is called a simplicial vertex of $G$ if $N_{G}[v]$ induces a clique in $G$. A perfect elimination ordering (PEO) of $G$ is an ordering of the vertices, say $\sigma=(v_{1},v_{2},\ldots,v_{n})$, such that $v_{i}$ is a simplicial vertex of $G_{i}=G[\{v_{i},v_{i+1},\ldots,v_{n}\}]$ for all $1\leq i\leq n$. The existence of PEO can characterize chordal graphs; that is, a graph $G$ is chordal if and only if $G$ has a PEO [fulkerson1965incidence].
Let $F$ be a family of sets. The intersection graph of $F$ is obtained by representing each set in $F$ as a vertex, with an edge between two vertices whenever the corresponding sets intersect. A graph $G$ is an interval graph if it arises as the intersection graph of a family of intervals on the real line. If, in addition, no interval properly contains another, then $G$ is called a proper interval graph.
For a chordal graph $G$, a perfect elimination ordering (PEO) $\alpha=(v_{1},v_{2},\ldots,v_{n})$ is called a bi-compatible elimination ordering (BCO) if the reverse ordering $\alpha^{-1}=(v_{n},v_{n-1},\ldots,v_{1})$ is also a PEO of $G$. A graph $G$ admits a BCO if and only if it is a proper interval graph [jamison1982elimination]. The following observation concerns the structure of BCOs in proper interval graphs.
Observation 2 ([panda2003linear]).
If $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$ is a BCO of a connected proper interval graph $G=(V,E)$, then the set $v_{i},v_{i+1},\ldots,v_{j}$} forms a clique in $G$ if $v_{i}v_{j}\in E$ for some $i<j$.
A vertex $v\in V$ is called a cut vertex of $G=(V,E)$ if removal of $v$ increases the number of components in $G$. A block is a maximal connected subgraph of $G$, without any cut vertex. A graph $G$ is called a block graph if every block of $G$ is a clique. Two blocks are called adjacent if they contain a common cut vertex of $G$. A block graph with one or more cut vertices contains at least two blocks, each of which contain exactly one cut vertex. Such blocks are called end blocks. A graph $G=(V,E)$ is said to be a split graph if $V$ can be partitioned into an independent set $S$ and a clique $K$.
The main contributions of this paper are summarized below:
Our contributions.
  • We design a polynomial-time algorithm to determine the edge open packing number of a proper interval graph.
  • We design a polynomial-time algorithm to determine the edge open packing number of a block graph.
  • We show that the edge open packing number can be computed in linear time for split graphs.
The rest of the paper is organized as follows. In Section 2, we design a polynomial-time algorithm to determine the edge open packing number of a proper interval graph. Sections 3 describes polynomial-time algorithms for block graphs. Section 4 deals with linear-time algorithm for the edge open packing number in split graphs. Finally, Section 5 concludes the article.
2 EOP set in proper interval graphs
In this section, we present a polynomial-time algorithm for computing a maximum edge open packing set in a proper interval graph. Let G be a connected proper interval graph with a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$. For each $v_{i}$, $i\in[n]$, let $l(v_{i})=\max\left\{\{i\}\cup\{r:v_{i}v_{r}\in E(G)\right$. and $\left.r>i\right\}\}$. Let $v_{k}=l(v_{1})$. An edge $e=uv$ saturates its endpoints $u$ and $v$. Given a set of edges $S$, let $V_{S}$ denote the set of all vertices saturated by the edges in $S$. To design the algorithm, we need the following versions of optimal EOP sets in a proper interval graph.
  • $\rho_{e}^{o}\left(G\right)=\max\left\{|S|:S\right$. is an EOP set in $\left.G\right\}$;
  • $\rho_{1}\left(G\right)=\max\left\{|S|:S\right$. is an EOP set in $G$, where $v_{1}v_{j}\in S$ for some $\left.1<j\leq k\right\}$;
    • $\rho^{\prime}\left(G\right)=\max\left\{|S|:S\right$. is an EOP set in $G$, where $v_{1}v_{j}\in S$ for some $1<j\leq k$ and $v_{j}v_{p}\notin S$, for all $\left.k<p\leq l(v_{j})\right\}$;
    • $\rho^{\prime\prime}\left(G\right)=\max\left\{|S|:S\right$. is an EOP set in $G$, where $v_{1}v_{j}\in S$ for some $1<j\leq k$ and $v_{j}v_{p}\in S$, for some $\left.k<p\leq l(v_{j})\right\}$;
      (The above two parameters are used to calculate the value of $\rho_{1}\left(G\right)$)
  • $\rho_{2}\left(G\right)=\max\left\{|S|:S\right$. is an EOP set in $G$, where $v_{1}v_{j}\notin S$, for all $\left.1<j\leq k\right\}$;
Let G be a connected proper interval graph with a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$ and $S$ be an arbitrary EOP set of $G$. Since the vertices from $A=\{v_{1},v_{2},\ldots,v_{k}\}$ form a clique in $G$, in $S$ at most one edge from the clique $G[A]$ can be present. Then one of the following possibilities occurs: $(i)$ $v_{1}v_{j}\in S$ for some $1<j\leq k$, or $(ii)$ $v_{1}$ is not incident with an edge from $S$, that is, $v_{1}\notin V_{S}$. Based on this, we have the following observation.
Observation 3.
Let G be a connected proper interval graph with a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$. Then $\rho_{e}^{o}\left(G\right)=\max\left\{\rho_{1}\left(G\right),\rho_{2}\left(G\right)\right\}$.
In the subsequent lemmas, we determine the values of the two parameters $\rho_{1}\left(G\right)$, and $\rho_{2}\left(G\right)$, respectively.
Lemma 4.
Let G be a connected proper interval graph with a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$. Then
  1. (i)
    $\rho^{\prime}\left(G\right)=1+\displaystyle\max_{1<j\leq k}\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$, where $\mathcal{H}_{j}=G[v_{l(v_{j})+1},v_{l(v_{j})+2},\ldots,v_{n}]$;
  2. (ii)
    $\rho^{\prime\prime}\left(G\right)=2+\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$, where $\mathcal{G}^{p}_{j}=G[v_{l(v_{p})+1},v_{l(v_{p})+2},\ldots,v_{n}]$.
Moreover, $\rho_{1}\left(G\right)=\max\left\{\rho^{\prime}\left(G\right),\rho^{\prime\prime}\left(G\right)\right\}$.
Proof.
Let $S$ be an EOP set of $G$ that satisfies the conditions from the definition of $\rho_{1}\left(G\right)$. That is $v_{1}v_{j}\in S$ for some $1<j\leq k$, where $v_{k}=l(v_{1})$. As $G$ is a proper interval graph and $\sigma$ is a BCO, by the Observation 2, $\{v_{1},v_{2},\ldots,v_{k}\}$ induced a clique in $G$. For a fixed $j$, since $S$ is an EOP set and contains the edge $v_{1}v_{j}$, no other edges having an end vertex from $\{v_{1},v_{2},\ldots,v_{j-1},v_{j+1},\ldots,v_{k}\}$ can be in $S$. As $G[\{v_{1},v_{2},\ldots,v_{k}\}]$ is a clique in $G$, no edges of the form $v_{j}v_{r}$, for $2<r\leq k$ and $r\neq j$, can be in $S$. Moreover, at most one edge of the form $v_{j}v_{p}$, for $k<p\leq l(v_{j})$, can be in $S$ as $G[\{v_{j},v_{j+1},\ldots,v_{l(v_{j})}\}]$ form a clique in $G$. Based on this, we have the following two cases:
Case 1.
$v_{1}v_{j}\in S$ for some $1<j\leq k$ and $v_{j}v_{p}\notin S$, for all $k<p\leq l(v_{j})$
Let us assume $\rho^{\prime}\left(G\right)=|S|$. Now fix a $j$, and consider $S_{j}=S\setminus\{v_{1}v_{j}\}$ for $1<j\leq k$. By the condition, we have $S_{j}$ is an EOP set of $\mathcal{H}_{j}$, where $\mathcal{H}_{j}=G[v_{l(v_{j})+1},v_{l(v_{j})+2},\ldots,v_{n}]$. Thus, $|S_{j}|\leq\rho_{e}^{o}(\mathcal{H}_{j})$. Therefore, by the definition of $\rho^{\prime}\left(G\right)$, we have $\rho^{\prime}\left(G\right)=|S|=1+|S_{j}|\leq 1+\displaystyle\max_{1<j\leq k}{|S_{j}|}\leq 1+\displaystyle\max_{1<j\leq k}\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$.
On the other hand, let $S_{j}$ be a $\rho_{e}^{o}(\mathcal{H}_{j})$-set. Furthermore, let $S^{\prime}=S_{j}\cup\{v_{1}v_{j}\}$. Clearly, $S^{\prime}$ is an EOP set of $G$ such that $v_{1}v_{j}\in S^{\prime}$ for some $1<j\leq k$ and $v_{j}v_{p}\notin S^{\prime}$ for all $k<p\leq l(v_{j})$, and this is true for any $j$. Let us consider $j$, such that $|S_{j}|=\displaystyle\max_{1<j\leq k}\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$. Therefore, $\rho^{\prime}\left(G\right)\geq|S^{\prime}|=1+|S_{j}|=1+\displaystyle\max_{1<j\leq k}\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$. Hence, we have $\rho^{\prime}\left(G\right)=1+\displaystyle\max_{1<j\leq k}\rho_{e}^{o}\left(\mathcal{H}_{j}\right)$ (i).
Case 2.
$v_{1}v_{j}\in S$ for some $1<j\leq k$ and $v_{j}v_{p}\in S$ for some $k<p\leq l(v_{j})$
Let us assume $\rho^{\prime\prime}\left(G\right)=|S|$. Now fix a $j$ and $p$, and consider $S_{j}^{p}=S\setminus\{v_{1}v_{j},v_{j}v_{p}\}$ for $1<j\leq k$ and $k<p\leq l(v_{j})$. Since $S$ is an EOP set of $G$ and contains $v_{1}v_{j},v_{j}v_{p}$, no edges having an end vertex from $\{v_{1},v_{2},\ldots,v_{k},v_{k+1},\ldots v_{p},v_{p+1},\ldots,v_{l(v_{p})}\}$ apart from $v_{1}v_{j},v_{j}v_{p}$ can be in $S$. Therefore, $S_{j}^{p}$ is an EOP set of $\mathcal{G}^{p}_{j}$, where $\mathcal{G}^{p}_{j}=G[v_{l(v_{p})+1},v_{l(v_{p})+2},\ldots,v_{n}]$. Thus, $|S_{j}^{p}|\leq\rho_{e}^{o}(\mathcal{G}^{p}_{j})$. Hence, by the definition of $\rho^{\prime\prime}\left(G\right)$, we have $\rho^{\prime\prime}\left(G\right)=|S|=2+|S_{j}^{p}|\leq 2+\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}{|S_{j}^{p}|}\leq 2+\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$.
On the other hand, let $S_{j}^{p}$ be a $\rho_{e}^{o}(\mathcal{G}^{p}_{j})$-set. Furthermore, let $S^{\prime}=S_{j}\cup\{v_{1}v_{j},v_{j}v_{p}\}$. Clearly, $S^{\prime}$ is an EOP set of $G$ such that $v_{1}v_{j}\in S^{\prime}$ for some $1<j\leq k$ and $v_{j}v_{p}\in S^{\prime}$ for some $k<p\leq l(v_{j})$, and this is true for any $j$ and $p$. Let us consider $j$ and $p$, such that $|S_{j}^{p}|=\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$. Therefore, $\rho^{\prime\prime}\left(G\right)\geq|S^{\prime}|=2+|S_{j}^{p}|=2+\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$. Hence, we have $\rho^{\prime\prime}\left(G\right)=2+\displaystyle\max_{\begin{subarray}{c}1<j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}\left(\mathcal{G}^{p}_{j}\right)$ (ii).
From the definition of $\rho_{1}\left(G\right),\rho^{\prime}\left(G\right)$ and $\rho^{\prime\prime}\left(G\right)$, it follows that $\rho_{1}\left(G\right)=\max\left\{\rho^{\prime}\left(G\right),\rho^{\prime\prime}\left(G\right)\right\}$. ∎
Lemma 5.
Let G be a connected proper interval graph with a BCO $\sigma=\{v_{1},v_{2},\ldots,v_{n}\}$. Then $\rho_{2}\left(G\right)=\rho_{e}^{o}\left(\mathcal{G}^{\prime}\right)$, where $\mathcal{G}^{\prime}=G[v_{2},v_{3},\ldots,v_{n}]$.
Proof.
Let $S$ be an EOP set of $G$ that satisfies the conditions from the definition of $\rho_{2}\left(G\right)$. That is $v_{1}\notin V_{S}$, where $V_{S}$ is the set of all vertices saturated by the edges in $S$. Since $v_{1}\notin V_{S}$, no edges of the form $v_{1}v_{j}$ can be in $S$. Therefore, $S$ is an EOP set of $\mathcal{G}^{\prime}=G[v_{2},v_{3},\ldots,v_{n}]$. However, any EOP set of $\mathcal{G}^{\prime}$ is also an EOP set of $G$, where $v_{1}$ cannot be an end vertex of any edge. Therefore, we have $\rho_{2}\left(G\right)=\rho_{e}^{o}\left(\mathcal{G}^{\prime}\right)$. ∎
Using the results discussed above, we propose an efficient algorithm for finding $\rho_{e}^{o}(G)$ for a connected proper interval graph $G$ in the algorithm MaxEOP_PIG(G)1. Also note that the algorithm MaxEOP_PIG(G)1 runs in polynomial time.
Input: A connected proper interval graph $G=(V,E)$ with $n$ vertices.
Output: $\rho_{e}^{o}(G)$, the edge open packing number of $G$.
1
2Compute a BCO $\sigma=(v_{1},v_{2},\dots,v_{n})$ of $G$;
3 Let $k\leftarrow l(v_{1})$;
4
5for $j\leftarrow 2$ to $k$ do
6   let $\mathcal{H}_{j}\leftarrow G[v_{\,l(v_{j})+1},\dots,v_{n}]$;
 $\rho_{e}^{o}(\mathcal{H}_{j})\leftarrow\,$MaxEOP_PIG($\mathcal{H}_{j}$);
 // recursive call
7 if $l(v_{j})\neq k$ then
8    for $p\leftarrow k+1$ to $l(v_{j})$ do
9         let $\mathcal{G}^{p}_{j}\leftarrow G[v_{\,l(v_{p})+1},\dots,v_{n}]$;
       $\rho_{e}^{o}(\mathcal{G}^{p}_{j})\leftarrow\,$MaxEOP_PIG($\mathcal{G}^{p}_{j}$);
       // recursive call
10
11
12
13
[…]
15 $\rho^{\prime\prime}(G)\leftarrow 2+\displaystyle\max_{\begin{subarray}{c}2\leq j\leq k\\ k<p\leq l(v_{j})\end{subarray}}\rho_{e}^{o}(\mathcal{G}^{p}_{j})$;
16 $\rho_{1}(G)\leftarrow\max\{\rho^{\prime}(G),\rho^{\prime\prime}(G)\}$;
17
18let $\mathcal{G}^{\prime}\leftarrow G[v_{2},v_{3},\dots,v_{n}]$;
19 $\rho_{2}(G)\leftarrow\rho_{e}^{o}(\mathcal{G}^{\prime})$;
20
[…]
22
23return $\rho_{e}^{o}(G)$;
Algorithm 1 MaxEOP_PIG($G$)
Theorem 6.
Maximum edge open packing can be solved in polynomial time for proper interval graphs.
3 EOP set in block graphs
In this section, we design a polynomial-time algorithm for finding the edge open packing number of a given block graph. We start with some pertinent definitions related to block graphs. Let $G$ be a block graph with $l$ blocks, $B_{1},B_{2},\ldots,B_{l}$, and $k$ cut vertices, $v_{1},v_{2},\ldots,v_{k}$. The cut-tree of $G$, denoted by $T_{G}=(V^{\prime},E^{\prime})$, is defined as $V^{\prime}=\{B_{1},B_{2},\ldots,B_{l},v_{1},v_{2},\ldots,v_{k}\}$ and $E^{\prime}=\{B_{i}v_{j}|v_{j}\in B_{i},1\leq i\leq l,1\leq j\leq k\}$. An example of a block graph and its cut-tree is illustrated in Figure 1. A vertex of $T_{G}$ corresponding to a block of $G$ is called a block vertex. The cut-tree $T_{G}$ of block graph $G$ can be made a rooted tree by selecting any block or cut vertex as the root. The set of all blocks of a given block graph $G=(V,E)$ can be found in $O(|V|+|E|)$ time by performing a depth-first search (DFS). For each cut vertex $v$ of $G$, the set of all blocks containing $v$ can also be found in $O(|V|+|E|)$ time. Hence the rooted cut-tree of a block graph can be constructed in $O(|V|+|E|)$ time [AhoHopcroftUllman].
Let $V(B)$ be the set of vertices of $B$ in $G$. For a cut vertex $r$, let $T_{G}^{r}$ be the rooted cut-tree of the block graph $G$, and let $r$ be the root of this rooted cut-tree. If $v$ is a non-cut vertex that belongs to the block $B_{j}$, then we define the graph $G_{v}$ as the connected component of the graph obtained from $G$ by deleting all the vertices $(V(B_{j})\setminus\{v\})$ and containing $v$. On the other hand, if $v$ is a cut vertex of $G$ and $B_{j}$ is the block corresponding to the parent block vertex of $v$ in $T_{G}^{r}$, then we define the graph $G_{v}$ as the connected component of the graph obtained from $G$ by deleting all the vertices $(V(B_{j})\setminus\{v\})$ and containing $v$. Note that when $v$ is a non-cut vertex, $G_{v}$ is just the single vertex $v$ and $G_{r}$ is isomorphic to $G$ with $r$ being the root of the cut-tree. For every cut vertex $v\in V(G)$, let $CC(v)$ denote the set of children of $v$ in $T_{G}^{r}$, and for a block vertex corresponding to a block $B$, let $CB(B)$ denote the children of $B$ in $T_{G}^{r}$. Let $v$ be a cut vertex and $G_{v}$ be the graph defined earlier. Then for a block $B\in CC(v)$, we define the graph $G_{v}^{B}$ as the connected component of $G_{v}\setminus\{v\}$ and part of block $B$. The graphs $G_{v}$ and $G_{v}^{B}$ is illustrated in Figure 1.
Figure 1: On the right is a block graph $G$, in the middle is its cut-tree $T_{G}^{v_{1}}$, and on the left side are the graphs $G_{v}$ and $G_{v}^{B}$, respectively
To design the algorithm, we need the following versions of optimal EOP sets in a block graph with respect to rooted cut-tree $T_{G}^{r}$.
  • $\rho_{e}^{o}\left(G_{v}\right)=\max\left\{|S|:S\right$. is an EOP set in $\left.G_{v}\right\}$;
  • $\rho^{c}\left(G_{v}\right)=\max\left\{|S|:S\right$. is an EOP set in $G_{v}$, where $v$ is the center of a star in $\left.G_{v}[S]\right\}$;
  • $\rho^{l}\left(G_{v}\right)=\max\left\{|S|:S\right$. is an EOP set in $G_{v}$, where $v$ is a leaf of a star in $\left.G_{v}[S]\right\}$;
  • $\rho^{\prime}\left(G_{v}\right)=\max\left\{|S|:S\right$. is an EOP set in $G_{v}$, where $v$ is not in $\left.V(G_{v}[S])\right\}$;
  • $\rho^{\prime\prime}\left(G_{v}\right)=\max\left\{|S|:S\right$. is an EOP set in $G_{v}$, where $\left.N[v]\cap V(G_{v}[S])=\emptyset\right\}$;
Let $G$ be a block graph and $T_{G}^{v}$ be its rooted cut-tree. Now if we take an arbitrary EOP set $S$ of $G_{v}$, we have three possibilities: $(a)$ $v$ is the center of a star from $G_{v}[S]$, $(b)$ $v$ is a leaf of a star from $G_{v}[S]$, or $(c)$ $v$ is not incident with an edge from $S$. If $v$ belongs to a $K_{1,1}$-component of $G_{v}[S]$, then we may consider $v$ either as the leaf or as the centre of the corresponding star $K_{1,1}$. Based on this, we have the following observation.
Observation 7.
Let $G$ be a connected block graph and $T_{G}^{v}$ be its rooted cut-tree. Then $\rho_{e}^{o}\left(G\right)=\rho_{e}^{o}\left(G_{v}\right)=\max\left\{\rho^{c}\left(G_{v}\right),\rho^{l}\left(G_{v}\right),\rho^{\prime}\left(G_{v}\right)\right\}$.
Note that for a non-cut vertex $v$, $G_{v}$ is a single-vertex graph, so $\rho^{c}\left(G_{v}\right)=\rho^{l}\left(G_{v}\right)=\rho^{\prime}\left(G_{v}\right)=\rho^{\prime\prime}\left(G_{v}\right)=0$. Now we want to calculate $\rho^{c}\left(G_{v}\right),\rho^{l}\left(G_{v}\right),\rho^{\prime}\left(G_{v}\right)$, and $\rho^{\prime\prime}\left(G_{v}\right)$ for a cut vertex $v$ in $G$.
For calculating $\rho^{c}\left(G_{v}\right)$, we need to define two types of children from $CC(v)$. Note that when $v$ is a centre of a star in an EOP set $S$ of $G_{v}$, there must be a block $B\in CC(v)$ such that $|E(G[B])\cap S|\geq 1$. Since every block is a complete graph, any EOP set of $G_{v}$ contains at most one edge from a block $B$. Hence, in our case $|E(G[B])\cap S|=1$ for some block $B$ in $CC(v)$. Therefore, for other blocks from $CC(v)$, either we have an edge in $S$ or no edges. Now we define two parameters corresponding to each block $B$ in $CC(v)$.
  1. (i)
    $\mathcal{A}_{B}=\displaystyle\max_{u\in V(B)\setminus\{v\}}\{1+\rho^{\prime\prime}\left(G_{u}\right)+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}$ if $S$ contains an edge from $B$
  2. (ii)
    $\mathcal{\overline{A}}_{B}=\displaystyle\sum_{w\in CB(B)}\rho^{\prime}\left(G_{w}\right)$ if $S$ does contains an edge from $B$
Note that the parameters $\mathcal{A}_{B}$ and $\mathcal{\overline{A}}_{B}$ are the edge open packing number of the block graph corresponding to the branch of $T_{G}^{v}$ containing $B$, when the maximum EOP set contains an edge from $B$ and no edges from $B$, respectively, subject to the condition that this maximum EOP set is contained in an EOP set $S$ of $G_{v}$, such that $v$ is a centre of a star in $G_{v}[S]$.
Now we say a block $B\in CC(v)$ is a Type 1 if $\mathcal{A}_{B}\geq\mathcal{\overline{A}}_{B}$, otherwise it is Type 2. Now we have the following lemma to calculate the value $\rho^{c}\left(G_{v}\right)$.
Lemma 8.
Let $G$ be a connected block graph with rooted cut-tree $T_{G}^{r}$. Then for any cut vertex $v$ in $G$,
[…]
Proof.
Note that since $G_{v}$ is a block graph and each block is a complete graph, form each block at most one edge can be in an EOP set of $G$. Let $S$ be an EOP set of $G_{v}$, which satisfies the conditions from the definition of $\rho^{c}\left(G_{v}\right)$. That is, $v$ is the centre of a star $\mathcal{S}$ in $G_{v}[S]$. Therefore, there exists a block $B\in CC(v)$, such that we have a vertex in $V(B)\setminus\{v\}$, which is a leaf of the star $\mathcal{S}$. Let $X\subseteq CC(v)$ such that from each block in $X$, there is a vertex $u$ such that $u$ is a leaf of the star $\mathcal{S}$. Also, let $X^{\prime}=CC(v)\setminus X$, it may me possible that $X^{\prime}=\emptyset$. For a $B\in CC(v)$ and for each $u\in CB(B)$, let $S_{u}^{B}=S\cap E\left(G_{u}\right)$. Note that
$|S|=\left|X\right|+\sum_{B\in CC(v)}\sum_{u\in CB(B)}\left|S_{u}^{B}\right|.$
Let $B\in X$ and $u\in V(B)\setminus\{v\}$ such that $u$ is a leaf of the star $\mathcal{S}$ in $G_{v}[S]$. As $v$ is the centre of the star $\mathcal{S}$ in $G_{v}[S]$ and since $uv\in S$, neither $u$ nor any vertex adjacent to $u$ in $G_{u}$ can be incident with an edge in $S_{u}^{B}$. So we have $\left|S_{u}^{B}\right|\leq\rho^{\prime\prime}\left(G_{u}\right)$. Let $w\in CB(B)\setminus\{u\}$. As $uv\in S$, $w$ cannot be incident with an edge in $S_{w}^{B}$, but its adjacent vertices in $G_{w}$ can be incident with an edge in $S_{w}^{B}$. Therefore, $\left|S_{w}^{B}\right|\leq\rho^{\prime}\left(G_{w}\right)$ for $w\in CB(B)\setminus\{u\}$. Now we assume the other situation when $B\in X^{\prime}$. In this case for all $\in CB(B), uu$ cannot be incident with an edge in $S_{u}$, but its adjacent vertices in $G_{u}$ can be incident with an edge in $S_{u}^{B}$. Therefore, $\left|S_{u}^{B}\right|\leq\rho^{\prime}\left(G_{u}\right)$ for all $u\in CB(B)$ when $B\in X^{\prime}$.
First, assume that there exists a block $B$ in $CC(v)$ of Type 1. Now, if $B\in X^{\prime}$ is of Type 1, then $\sum_{u\in CB(B)}\left|S_{u}^{B}\right|\leq\displaystyle\sum_{u\in CB(B)}\rho^{\prime}\left(G_{u}\right)=\mathcal{\overline{A}}_{B}\leq\mathcal{A}_{B}$. Similarly, if $B\in X$ is of Type 2, then $\sum_{u\in CB(B)}\left|S_{u}^{B}\right|\leq\displaystyle\max_{u\in V(B)\setminus\{v\}}\{\rho^{\prime\prime}\left(G_{u}\right)+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}=\mathcal{A}_{B}-1\leq\mathcal{\overline{A}}_{B}-2$. Let us denote $L_{B}=\displaystyle\sum_{u\in CB(B)}\left|S_{u}^{B}\right|$. Thus,
[…]
$\leq\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 1}}\end{subarray}}\mathcal{A}_{B}+\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}$
Now assume all the children of $CC(v)$ are of Type 2. Then,
$|S|$ $=\left|X\right|+\sum_{\begin{subarray}{c}B\in X\\ B\text{ {Type 2}}\end{subarray}}L_{B}+\sum_{\begin{subarray}{c}B\in X^{\prime}\\ B\text{ {Type 2}}\end{subarray}}L_{B}$
$\leq\left|X\right|+\sum_{\begin{subarray}{c}B\in X\\ B\text{ {Type 2}}\end{subarray}}(\mathcal{A}_{B}-1)+\sum_{\begin{subarray}{c}B\in X^{\prime}\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}$
$=\sum_{\begin{subarray}{c}B\in X\\ B\text{ {Type 2}}\end{subarray}}(\mathcal{A}_{B}-\mathcal{\overline{A}}_{B}+\mathcal{\overline{A}}_{B})+\sum_{\begin{subarray}{c}B\in X^{\prime}\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}$
$=\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}+\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 2}}\end{subarray}}(\mathcal{A}_{B}-\mathcal{\overline{A}}_{B})$
$\leq\sum_{B\in CC(v)}\mathcal{\overline{A}}_{B}+\max_{B\in CC(v)}\{\mathcal{A}_{B}-\mathcal{\overline{A}}_{B}\}[\textbf{ As }\mathcal{A}_{B}<\mathcal{\overline{A}}_{B}].$
Consequently, it follows that $\rho^{c}(G_{v})$ cannot exceed the expression given on the right-hand side of the lemma.
We also look at two situations for the reverse inequality. Let there be an element $B\in CC(v)$ such that $\mathcal{A}_{B}\geq\mathcal{\overline{A}}_{B}$; that is, $B$ is a Type 1 child of $v$ in $T^{r}_{v}$. Let $D$ be an EOP set of $G_{v}$, where $v$ is the centre of a star in $G_{v}[D]$, such that $vu\in D$ for some $u\in V(B)\setminus\{v\}$ if and only if $B$ is of Type 1 while $G_{u}[D]$ for some $u\in CB(B)$ depends on the type of $B\in CC(v)$. That is, if $B$ is of Type 1, then let $E\left(G_{u}[B]\right)$ be a $\rho^{\prime\prime}\left(G_{u}\right)$-set for exactly one $u\in V(B)\setminus\{v\}$ and for the other vertex $w\in CB(B)\setminus\{u\}$, let $E\left(G_{w}[B]\right)$ be a $\rho^{\prime}\left(G_{w}\right)$-set. On the other hand, if $B$ is a Type 2, then let $E\left(G_{w}[B]\right)$ be a $\rho^{\prime}\left(G_{w}\right)$-set for every $w\in CB(B)$. Consequently, the inequality
$\rho^{c}\left(G_{v}\right)$ $\geq\left|D\right|=\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 1}}\end{subarray}}(\max_{u\in V(B)\setminus\{v\}}\{1+\rho^{\prime\prime}\left(G_{u}\right)+\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\})+\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 2}}\end{subarray}}\sum_{w\in CB(B)}\rho^{\prime}\left(G_{w}\right)$
$=\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 1}}\end{subarray}}\mathcal{A}_{B}+\sum_{\begin{subarray}{c}B\in CC(v)\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}$
is clearly established. Now assume all children of $v$ in $T^{r}_{v}$ are of Type 2. Let $B^{*}$ be a block in $CC(v)$ such that $(\mathcal{A}_{B^{*}}-1)-\mathcal{\overline{A}}_{B^{*}}$ is the largest possible among all vertices in $CC(v)$. Letting $\mathcal{D^{*}}$ be an EOP set in $G_{v}$, in which $v$ is the centre of the star $\mathcal{S}$ containing only $v$ and $u$, for some $u\in V(B^{*})\setminus\{v\}$, we deduce
$\rho^{c}\left(G_{v}\right)$ $\geq\left|D^{*}\right|\geq\sum_{\begin{subarray}{c}B\in CC(v)\setminus\{B^{*}\}\end{subarray}}\sum_{w\in CB(B)}\rho^{\prime}\left(G_{w}\right)+\mathcal{A}_{B^{*}}$
$=\sum_{\begin{subarray}{c}B\in CC(v)\setminus\{B^{*}\}\end{subarray}}\mathcal{\overline{A}}_{B}+\mathcal{A}_{B^{*}}$
$=\sum_{\begin{subarray}{c}B\in CC(v)\end{subarray}}\mathcal{\overline{A}}_{B}+\mathcal{A}_{B^{*}}-\mathcal{\overline{A}}_{B^{*}}$
$=\sum_{\begin{subarray}{c}B\in CC(v)\end{subarray}}\mathcal{\overline{A}}_{B}+\max_{B\in CC(v)}\{\mathcal{A}_{B}-\mathcal{\overline{A}}_{B}\}.$
This leads to the required equality. The proof is complete. ∎
Lemma 9.
Let $G$ be a connected block graph with rooted cut-tree $T_{G}^{r}$. Then for any cut vertex $v$ in $G$,
[…]
  1. (i)
    $\mathcal{M}_{B}=\displaystyle\max_{u\in V(B)\setminus\{v\}}\{\max\{\rho^{c}\left(G_{u}\right),\rho^{\prime\prime}\left(G_{u}\right)\}+1+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}$ and
  2. (ii)
[…]
Proof.
Let $S$ be an EOP set of $G_{v}$, which satisfies the conditions from the definition of $\rho^{l}\left(G_{v}\right)$. That is, $v$ is a leaf of a star $\mathcal{S}$ in $G_{v}[S]$. Therefore, there exists exactly one block $B\in CC(v)$, such that for exactly one vertex $\in V(B)\setminus\{v\}, uu$ is the centre of the star $\mathcal{S}$. Since $G_{v}$ is a block graph and each block is a complete graph, from each block at most one edge can be in an EOP set of $G_{v}$. Clearly, none of the edges from block $B^{\prime}\in CC(v)\setminus\{B\}$ belongs to $S$. Moreover, since $S$ is an EOP set of $G_{v}$, no vertex $w\in CB(B^{\prime})$ for $B^{\prime}\in CC(v)\setminus\{B\}$ is incident with an edge from $S$ and no vertex $w\in CB(B)\setminus\{u\}$ is incident with an edge from $S$. For any block $B\in CC(v)$ and for each $u\in CB(B)$, let $S_{u}^{B}=S\cap E\left(G_{u}\right)$. If $B^{\prime}\neq B$, then $S_{w}^{B^{\prime}}$ is an EOP set of $G_{w}$ such that $w\notin V\left(G_{w}\left[S_{w}^{B^{\prime}}\right]\right)$. Therefore, $\sum_{w\in CB(B^{\prime})}\left|S_{w}^{B^{\prime}}\right|\leq\displaystyle\sum_{w\in CB(B^{\prime})}\rho^{\prime}\left(G_{w}\right)$.
On the other hand, since $u$ is the centre of the star $\mathcal{S}$ in $G_{v}[S]$, there are two possibilities for the edges from $S_{u}^{B}$ depending on the order of $\mathcal{S}$. If $\mathcal{S}$ is a $K_{1,1}$, then none of the vertices in $N[u]\cap V\left(G_{u}\right)$ can be in $G_{u}\left[S_{u}^{B}\right]$, and so $\left|S_{u}^{B}\right|\leq\rho^{\prime\prime}\left(G_{u}\right)$. If $|V(\mathcal{S})|\geq 3$, then some of the edges between $u$ and vertices in $G_{u}$ are in $S_{u}^{B}$. Hence, $\left|S_{u}^{B}\right|\leq\rho^{c}\left(G_{u}\right)$. But in both cases, for the other vertex $w\in CB(B)\setminus\{u\}$, $S_{w}^{B}$ is an EOP set of $G_{w}$ such that $w\notin V\left(G_{w}\left[S_{w}^{B}\right]\right)$. Now consider all cases; we have $\sum_{w\in CB(B)}\left|S_{w}^{B}\right|\leq\displaystyle\max_{u\in V(B)\setminus\{v\}}\{\max\{\rho^{c}\left(G_{u}\right),\rho^{\prime\prime}\left(G_{u}\right)\}+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}$. Finally, combining for all block $\tilde{B}\in CC(v)$, we get
$|S|=$ $\sum_{w\in CB(B)}\left|S_{w}^{B}\right|+1+\sum_{B^{\prime}\in CC(v)\setminus\{B\}}\sum_{w\in CB(B^{\prime})}\left|S_{w}^{B^{\prime}}\right|$
$\leq\max_{u\in V(B)\setminus\{v\}}\{\max\{\rho^{c}\left(G_{u}\right),\rho^{\prime\prime}\left(G_{u}\right)\}+1+\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}+$
$\sum_{B^{\prime}\in CC(v)\setminus\{B\}}\sum_{w\in CB(B^{\prime})}\rho^{\prime}\left(G_{w}\right)$
[…]
Conversely, the construction outlined in the preceding paragraph results in an EOP set $S$ of $G_{v}$ with the specified cardinality. In this construction, $v$ represents a leaf of a star $\mathcal{S}$ within $G_{v}[S]$. For the corresponding one-block vertex $B\in CC(v)$ and a vertex $\in V(B)\setminus\{v\}, uu$ serves as the centre of $\mathcal{S}$, and the intersection $S\cap E\left(G_{u}\right)$ constitutes either a $\rho^{c}\left(G_{u}\right)$-set or a $\rho^{\prime\prime}\left(G_{u}\right)$-set, contingent upon which set is larger. Furthermore, in any other sets $S\cap E\left(G_{w}\right)$, where $w\in CB(B)\setminus\{u\}$, we can select a $\rho^{\prime}\left(G_{w}\right)$-set. Moreover, in every other set $S\cap E\left(G_{w}\right)$, where $w\in CB(B^{\prime})$ and $B^{\prime}\in CC(v)\setminus\{B\}$, we can select a $\rho^{\prime}\left(G_{w}\right)$-set. For a certain $B$, we first determine the maximum across all $u\in V(B)\setminus\{v\}$ and subsequently identify the largest feasible set $S$ across all $B\in CC(v)$, therefore confirming that the asserted value for $|S|$ is attained. ∎
Lemma 10.
Let $G$ be a connected block graph with rooted cut-tree $T_{G}^{r}$. Then for any cut vertex $v$ in $G$, $\rho^{\prime}\left(G_{v}\right)=\displaystyle\sum_{B\in CC(v)}\rho_{e}^{o}\left(G_{v}^{B}\right)$.
Proof.
Let $S$ be an EOP set of $G_{v}$ that fulfils the criteria outlined in the definition of $\rho^{\prime}\left(G_{v}\right)$. Thus, $v$ is not associated with an edge from $S$. For any $B\in CC(v)$, define $S_{B}=S\cap E(G_{u})$. It is important to observe that $S_{B}$ constitutes an EOP set of $G_{v}^{B}$, so $\left|S_{B}\right|\leq\rho_{e}^{o}\left(G_{v}^{B}\right)$. Consequently, $|S|=\displaystyle\sum_{B\in CC(v)}\left|S_{B}\right|\leq\displaystyle\sum_{B\in CC(v)}\rho_{e}^{o}\left(G_{v}^{B}\right)$. Conversely, by selecting a $\rho_{e}^{o}\left(G_{v}^{B}\right)$-set $S_{B}$ in $G_{v}^{B}$ for each $B\in CC(v)$, the sum of all these sets $S_{B}$ produces an EOP set of $G_{v}$ with a cardinality of $\sum_{B\in CC(v)}\rho_{e}^{o}\left(G_{v}^{B}\right)$, so concluding the proof of the lemma. ∎
Lemma 11.
Let $G$ be a connected block graph with rooted cut-tree $T_{G}^{r}$. Then for any cut vertex $v$ in $G$, $\rho^{\prime\prime}\left(G_{v}\right)=\displaystyle\sum_{B\in CC(v)}\sum_{u\in CB(B)}\rho^{\prime}\left(G_{u}\right)$.
Proof.
Let $S$ be an EOP set of $G_{v}$ that fulfils the criteria outlined in the definition of $\rho^{\prime\prime}\left(G_{v}\right)$. Thus, $N[v]$ is not associated with an edge from $S$. Since $G_{x}$ is a single vertex graph for any non-cut vertex $x\in V(B)$, we only consider the cut vertices of every block $B\in CC(v)$. For any block $B\in CC(v)$ and for each $u\in CB(B)$, let $S_{u}^{B}=S\cap E\left(G_{u}\right)$. Note that $S_{u}^{B}$ satisfies the conditions from the definition of $\rho^{\prime}\left(G_{u}\right)$, and so $\left|S_{u}^{B}\right|\leq\rho^{\prime}\left(G_{u}\right)$. We deduce that $|S|=\displaystyle\sum_{B\in CC(v)}\sum_{u\in CB(B)}\left|S_{u}^{B}\right|\leq\displaystyle\sum_{B\in CC(v)}\sum_{u\in CB(B)}\rho^{\prime}\left(G_{u}\right)$. Conversely, selecting the biggest EOP set $S_{u}^{B}$ in $G_{u}$ for each $u\in CB(B)$ and for every $B\in CC(v)$, ensuring that $u$ is not incident to an edge from $S_{u}^{B}$, the union $S^{*}$ of all these sets $S_{u}^{B}$ produces an EOP set of $G_{v}$ of cardinality $\sum_{B\in CC(v)}\sum_{u\in CB(B)}\rho^{\prime}\left(G_{u}\right)$. Clearly, $S^{*}$ constitutes an EOP set of $G_{v}$, such that $N[v]\cap V\left(G_{v}\right)[S^{*}]=\emptyset$, so concluding the proof of the lemma. ∎
Based on the above statements, we propose an efficient algorithm for finding $\rho_{e}^{o}(G)$ for a connected block graph $G$ in the algorithm MaxEOP_Block(G)2.
Input: A connected block graph $G=(V,E)$ with $n$ vertices.
Output: $\rho_{e}^{o}(G)$, the edge open packing number of $G$.
1
2Create a rooted cut-tree $T_{G}^{v}$, rooted at a cut vertex $v$;
3
4Let $\sigma=(v_{1},v_{2},\ldots,v_{k}=v)$ be the ordering of the cut vertices of $G$ such that $d_{T_{G}^{v}}(v,v_{1})\geq d_{T_{G}^{v}}(v,v_{2})\geq\ldots\geq d_{T_{G}^{v}}(v,v_{k})$;
5
6For a non-cut vertex $u$, $\rho_{e}^{o}(G_{u})=0$;
7
8for $j=1$ to $k$ do
9
10 $\rho^{\prime\prime}\left(G_{v_{i}}\right)\leftarrow\displaystyle\sum_{B\in CC(v_{i})}\sum_{u\in CB(B)}\rho^{\prime}\left(G_{u}\right)$;
11 $\rho^{\prime}\left(G_{v_{i}}\right)\leftarrow\displaystyle\sum_{B\in CC(v_{i})}\rho_{e}^{o}\left(G_{v_{i}}^{B}\right)$;
12
13 for $B\in CC(v_{i})$ do
14
15    $\mathcal{A}_{B}=\displaystyle\max_{u\in V(B)\setminus\{v\}}\{1+\rho^{\prime\prime}\left(G_{u}\right)+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}$;
16
17    $\mathcal{\overline{A}}_{B}=\displaystyle\sum_{w\in CB(B)}\rho^{\prime}\left(G_{w}\right)$;
18
19
20 if $\mathcal{A}_{B}\geq\mathcal{\overline{A}}_{B}$ then
21
22    $B\leftarrow$ Type 1;
23
24
25 else
26    $B\leftarrow$ Type 2;
27
28
29 if $\forall B\in CC(v_{i})[B=\text{ {Type 2}}]$ then
30
31    $\rho^{c}\left(G_{v_{i}}\right)\leftarrow\displaystyle\sum_{B\in CC(v_{i})}\mathcal{\overline{A}}_{B}+\displaystyle\max_{B\in CC(v_{i})}\{\mathcal{A}_{B}-\mathcal{\overline{A}}_{B}\}$;
32
33
34 else
35    $\rho^{c}\left(G_{v_{i}}\right)\leftarrow\displaystyle\sum_{\begin{subarray}{c}B\in CC(v_{i})\\ B\text{ {Type 1}}\end{subarray}}\mathcal{A}_{B}+\sum_{\begin{subarray}{c}B\in CC(v_{i})\\ B\text{ {Type 2}}\end{subarray}}\mathcal{\overline{A}}_{B}$;
36
37
38 for $B\in CC(v_{i})$ do
39
40    $\mathcal{M}_{B}=\displaystyle\max_{u\in V(B)\setminus\{v\}}\{\max\{\rho^{c}\left(G_{u}\right),\rho^{\prime\prime}\left(G_{u}\right)\}+1+\displaystyle\sum_{w\in CB(B)\setminus\{u\}}\rho^{\prime}\left(G_{w}\right)\}$;
41
[…]
43
44
[…]
46
47 $\rho_{e}^{o}\left(G\right)\leftarrow\rho_{e}^{o}\left(G_{v_{i}}\right)=\max\left\{\rho^{c}\left(G_{v_{i}}\right),\rho^{l}\left(G_{v}\right),\rho^{\prime}\left(G_{v_{i}}\right)\right\}$;
48
49
50return $\rho_{e}^{o}\left(G\right)$;
51
Algorithm 2 MaxEOP_Block(G)
Consequent to Observation 2 and Lemmas 8, 9, 10, and 11, we conclude that Algorithm 2 accurately computes the EOP number of a block graph $G$. Also note that the algorithm MaxEOP_Block(G)2 runs in polynomial time. Therefore, we have the following theorem.
Theorem 12.
Maximum edge open packing can be solved in polynomial time for block graphs.
4 EOP set in split graphs
In this section, we show that the edge open packing number of a split graph can be solved in linear time. Throughout this section, we assume that $G=(S\cup K,E)$ is a split graph, where $S$ and $K$ are independent sets and cliques of $G$. Furthermore, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. For all $x,y\in K$, we denote $l_{x}=|S\cap N(x)|,l_{y}=|S\cap N(y)|$, and $l_{xy}=|S\cap N(x)\cap N(y)|$. Then we have the following theorem.
Theorem 13.
Let $G=(S\cup K,E)$ be a split graph. Then $\rho_{e}^{o}\left(G\right)=\max\{\rho_{1}(G),\rho_{2}(G)\}$, where $\rho_{1}(G)=\displaystyle\max_{x,y\in K\\ x\neq y}\{l_{x}-l_{xy}+1,l_{y}-l_{xy}+1\}$ and $\rho_{2}(G)=\displaystyle\max_{x\in K}\{l_{x}\}$.
Proof.
Let us first assume $S=\emptyset$; clearly the graph is a complete graph, and we know that for a complete graph with more than one vertex $\rho_{e}^{o}\left(K_{n}\right)=1$. Note that in this case $l_{x}=l_{y}=l_{xy}=0$, and $\rho_{1}(G)=1$, $\rho_{2}(G)=0$. Therefore, $\rho_{e}^{o}\left(G\right)=1=\max\{\rho_{1}(G),\rho_{2}(G)\}$. Now we assume $S\neq\emptyset$. Let $D$ be an $\rho_{e}^{o}$-set of $G$. Since $K$ is a complete graph, $D$ can contain at most one edge from $E(G[K])$. First, assume $D$ contains an edge $e=xy$ of $E(G[K])$. As $D$ is an EOP set, no other edge from $E(G[K])$ can be in $D$. Then, $G[D]$ is a star centre at either $x$ or $y$. Also, note that if $z\in S\cap N(x)\cap N(y)$, then $z$ cannot be a leaf of the $G[D]$, as $xz,xy$ have a common edge $yz$. So, in this case $|D|=\rho_{1}(G)=\displaystyle\max_{x,y\in K\\ x\neq y}\{l_{x}-l_{xy}+1,l_{y}-l_{xy}+1\}$. For the other case, when $D$ does not contain any edge from $E(G[K])$, $G[D]$ is a star centred at any vertex from $K$. Hence, in this case $D=\rho_{2}(G)=\displaystyle\max_{x\in K}\{l_{x}\}$. Therefore, $\rho_{e}^{o}\left(G\right)=\max\{\rho_{1}(G),\rho_{2}(G)\}$. ∎
5 Conclusion
In this paper, we have partially solved the open question posed by Brešar and Samadi about edge open packing sets in chordal graphs. First, we have designed a polynomial-time algorithm to solve the maximum edge open packing set in a proper interval graph. Furthermore, we have presented a polynomial-time algorithm to solve the maximum edge open packing set in block graphs. Finally, we have shown that the problem of edge open packing number can be solved in linear time in split graphs. It would be fascinating to investigate the other open problems mentioned in [chelladurai2022edge] and [brevsar2024edge].